OpenStudy (anonymous):

Solve this inequality. l2x-3l(is greater than or equal to)1

OpenStudy (anonymous):

\[|2x-3|\geq1\] \[2x-3\geq1 \space \space or\space \space 2x-3\leq-1\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Wait..... Isn't the negative three supposed to be a plus?

OpenStudy (anonymous):

no, it should be negative You have to solve for x but I trust you can do it

OpenStudy (anonymous):

Yeah, I can, I just thought it was positive because anything negative in absolute value turns positive

OpenStudy (anonymous):

The absolute value is over the whole thing not just the 3 if it were \[2x+|-3|\geq1\] then -3 would turn positive

OpenStudy (anonymous):

oooooh, so if it's by itself then it turns pos., but if it includes other numbers then it stays the same?

OpenStudy (anonymous):

\[|-(2x-3)|=(2x-3)\]

OpenStudy (anonymous):

2x+3???????

OpenStudy (anonymous):

So when we say \[|2x-3|\geq1\] What we are really is saying is that \[(2x-3)\geq1\] and \[-(2x-3)\geq1\]

OpenStudy (anonymous):

\[-(2x-3)\geq1\] When we divide both side by -1 we switch the inequality sign \[(2x-3)\leq-1\]

OpenStudy (anonymous):

hmmmmm, thx anyway. I'm so confused

OpenStudy (anonymous):

So the answers for the original two would be...... x(is greater than or equal to)2 OR x(is greater than or equal to)1?????

OpenStudy (anonymous):

I don't get what that means so can you tell me yes or no please?

OpenStudy (anonymous):

\[x\geq2\]or \[x\leq1\]

OpenStudy (anonymous):

YAY! I GOT IT RIGHT! THANK YOU!